a cylindrical vessel contains liquid of volume v0 and area of cross section ao at ttmprature t1 . Find expression for rise in level of liquid.
At t2


Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

your question is titled "expansion of liquid", so I am ignoring expansion of vessel.

let coefficient of volume expansion of liquid be 'gamma',

volume=aoh

ao*h2=ao*h1(1+(gamma)(t2-t1))=vo(1+(gamma)*(t2-t1))